# Thresholded Lexicographic Ordered Multiobjective Reinforcement Learning

### v1

@@ Abstract
Lexicographic multi-objective problems, which impose a lexicographic importance order over the objectives, arise in many real-life scenarios. Existing Reinforcement Learning work directly addressing lexicographic tasks has been scarce. The few proposed approaches were all noted to be heuristics without theoretical guarantees as the Bellman equation is not applicable to them. Additionally, the practical applicability of these prior approaches also suffers from various issues such as not being able to reach the goal state. While some of these issues have been known before, in this work we investigate further shortcomings, and propose fixes for improving practical performance in many cases. We also present a policy optimization approach using our Lexicographic Projection Optimization (LPO) algorithm that has the potential to address these theoretical and practical concerns. Finally, we demonstrate our proposed algorithms on benchmark problems.
@@ 1 Introduction
Our present work is in the a priori setting. The most common specification method in this setting is linear scalarization which which requires the designer to assign weights to the objectives and take a weighted sum, thus making solutions comparable [2]. The main benefit of this technique is that it allows the use of many standard off the shelf algorithms as it preserves the additivity of the reward functions. However, expressing user preference with this technique requires significant domain knowledge and preliminary work in most scenarios [8]. While it can be the preferred method when the objectives can be expressed in comparable quantities, e.g. when all objectives have a monetary value, this is not the case most of the time. Usually, the objectives are expressed in incomparable quantities like money, time, and carbon emissions. Additionally, a composite utility combining the various objectives, and an approximation of that with linear scalarization limits us to a subset of the Pareto optimal set.
To address these drawbacks of linear scalarization, several other approaches have been proposed and studied. Nonlinear scalarization methods like Chebyshev [13] are more expressive and can capture all of the solutions in the Pareto optimal set, however, they do not address the user-friendliness requirement. In this paper, we will focus on an alternative specification method that overcomes both limitations of linear scalarization, named Thresholded Lexicographic Ordering (TLO) [3] [8]. In lexicographic ordering, the user determines an importance order for the objectives. The more important objectives dominate the lower order objectives. Given two candidate solutions, a lower ranked objective is used to rank the two solutions only if all higher order objectives have the same values. The thresholding part of the technique allows a more generalized definition for being the same w.r.t. an objective. The user provides a threshold for each objective except the last, and the objective values are clipped at the corresponding thresholds. This allows the user to specify values beyond which they are indifferent to the optimization of an objective. There is no threshold for the last objective as it is considered an unconstrained open-ended objective.
@@ 2 Related Work
[3] was one of the first papers that investigate RL in multi-objective tasks with preference ordering. It introduces TLQ as an RL algorithm to solve such problems. Vamplew et al. [21] showed that TLQ significantly outperforms Linear Scalarization (LS) when the Pareto front is globally concave or when most of the solutions lie on the concave parts. However, LS performs better when the rewards are not restricted to terminal states, because TLQ cannot account for the already received rewards. Later, Roijers et al. [15] generalized this analysis by comparing more approaches using a unifying framework. To our knowledge, [21] is the only previous work that explicitly discussed shortcomings of TLQ. However, we found that TLQ has other significant issues that occur even outside of the problematic cases they analyze.
Wray et al. [25] introduced Lexicographic MDP (LMDP) and the Lexicographic Value Iteration (LVI) algorithm. LMDPs define the thresholds as slack variables which determine how worse than the optimal value is still sufficient. While Wray et al. [25] proved the convergence to desired policy if slacks are chosen appropriately, such slacks are generally too tight to allow user preferences. This is also observed in Pineda et al. [14] which claimed that while ignoring these slack bounds negates the theoretical guarantees, the resulting algorithm still can be useful in practice.
@@
However, we discovered that these works on using a Q-learning variant with thresholded ordering perform very poorly in most cases due to non-Markovianity of the value function they try to learn. It is possible to bypass this issue by using policy gradient approaches as they do not require learning an optimal value function. In order to handle conflicting gradients, some modifications to the gradient descent algorithm are needed. Recent work on modified gradient descent algorithms came mostly from Multi Task Learning literature, which could be considered a multiobjective optimization problem [1] [17] [10] [12] [11]. While these papers use similar ideas with our work, their setting is different than ours as they do not have any explicit importance order; hence, not applicable to our setting. Uchibe and Doya [20] have the most similar setting to ours in gradient-based algorithms. It considers a set of constraints with an unconstrained objective. Then, the gradient of the unconstrained objective is projected onto the positive half-space of the active (violated) constraints and adds a correction step to improve the active constraints. When no valid projection is found, the most violated constraints are ignored until a valid projection exists. This is one of the main differences with our setting: As we have an explicit importance-order of the objectives, it is not acceptable to ignore a constraint without considering the importance order. Also, we project the gradients onto hypercones instead of hyperplanes, which is a hypercone with $\pi/2$ vertex angle. Thus, our algorithm allows varying degrees of conservative projections to prevent a decline in the constraints.
While there are many other recent works on Constrained Markov Decision Process (CMDPs) like [23, 6], their approaches are not applicable as an importance order over constraints is not allowed. Recently, using RL with lexicographic ordering began to attract attention from other communities as well. For example, Hahn et al. [4] use formal methods to construct single objective MDPs when all of the objectives are $\omega$-regular.
@@
Skalse et al. [18] propose both value-based and policy-based approaches for LMDPs. Their value-based approach is based on slacks like [8] and they require using very small slack values. This protects their approach from the issues with having relaxations by limiting their setting to strict lexicographic order. For policy-based methods, they use Lagrangian relaxation and their setting is again a strict lexicographic ordering, i.e. it does not allow treating values above a threshold equal. As we are primarily interested in solving tasks with thresholded objectives, this paper does not address our question.
@@ 4.1 Shortcomings of Prior TLQ Approaches
Reachability Objective: A reachability objective is represented by a non-zero reward in the terminal states of the MDP and zero rewards elsewhere. We will call an objective that has non-zero rewards in at least one non-terminal state a non-reachability objective.
@@
(I) TLQ successful case studies: Vamplew et al. [21] show an experiment where the constrained objective is reachability, unconstrained objective is terminating, and TLQ works. Li and Czarnecki [8] show this for a case study where the constrained objective is a non-terminating reachability constraint. Note that these are just demonstrations of empirical performance for case studies that fall into these categories and the papers do not make claims about general instances.
(II) TLQ does not work: Vamplew et al. [21] show that TLQ does not work when constrained objective is non-reachability. To the best of our knowledge, this is the only failure case that has been shown previously in the literature.
@@
In addition to previously identified failure case in Vamplew et al. [21], TLQ does not work when the constrained objective is a terminating reachability objective but the unconstrained one is non-terminating.
@@
Problems with Reachability Constraint. A common scenario in control tasks is having a primary objective to reach a goal state and a secondary objective that evaluates the path taken to there. Formally, this is a scenario where the primary objective is a reachability objective. However, TLQ either needs to ignore the secondary objective or fails to satisfy the primary objective in such cases. All of the thresholding methods above fail to guarantee to reach the goal in this setting when used threshold/slack values are uniform throughout state space. The maze in Figure 1 can be used to observe this phenomenon. Assume that our primary objective is to reach $G$ and we encode this with a reward function that is 00 everywhere but $G$ where it is $R$. And our secondary objective is to avoid the bad tiles. A Pareto optimal policy in this maze could be indicated by the state trajectory $(1,0)\to(2,0)\to(2,1)\to(2,2)\to(1,2)$. However, this is unattainable by TLQ.
Since the reward is given only in the goal state, $\hat{Q}^{\star}_{1}$ can be equal to $\tau_{1}$ only for the state-action pairs that lead to the goal state. If no discounting was used, all actions would have the value $\tau_{1}$, hence the agent would not need to reach the goal state. We believe the reason why Vamplew et al. [21] have not observed this issue in their experiments with undiscounted ($\gamma=1$) Deep Sea Treasure (DST) is due to their objectives. The secondary objective of DST, minimizing the time steps before terminal state, is a terminating objective which pushes the agent to actually reach a goal state. Absolute Slacking also similarly leads to contradicting requirements when the trajectory described above is aimed but we leave this case to the reader due to space constraints.
@@ 5 Policy Gradient Approach for TLO
The proof and the exact lower bounds on $F_{j}(\boldsymbol{x}^{(l+1)})-F_{j}(\boldsymbol{x}^{(l)})$ in the proposition are in the supplementary material.
@@
Experiment. As a benchmark for the Lexicographic Projection Algorithm, we used $F_{1}(x,y)=-4x^{2}+-y^{2}+xy$ and $F_{2}(x,y)=-(x-1)^{2}-(y-0.5)^{2}$ which are taken from [27]. We modified $F_{2}$ slightly for better visualization and multiplied both functions with $-1$ to convert this to a maximization problem. We set the threshold for $F_{1}$ to $-0.5$. The behavior of our cone algorithm without using active constraints heuristic on this problem with $\tau=(-0.5)$ and $\Delta=\frac{\pi}{90}$ can be seen in Figure 2.
@@ 6 Experiments
We evaluate the performance of the Lexicographic REINFORCE algorithm on two Maze problems and a many-objective benchmark from the literature. In both experiments, we use a two layer neural network for policy function.
@@
Experiment 1: Non-reachability. The primary objective is non-reachability, i.e. it takes non-zero values in some non-terminal states. For this, we flip our objectives from the previous maze setting and define our primary objective as minimizing the cost incurred from the bad tiles. $HH$s give $-5$ reward and $hh$s give $-4$ reward. $+1$ reward is given in the terminal state to extend the primary objective to have rewards in both terminal and non-terminal states. The secondary objective is to minimize the time taken to the terminal state. We formalize this by defining our secondary reward function as 00 in terminal state and $-1$ everywhere else. We use the Maze in Figure 3 for this experiment. Note that this is the setting that Vamplew et al. [21] has found that TLQ fails. However, our experiments show that Lexicographic REINFORCE can solve this setting too.
Figure 4: Satisfaction rates for a single successful seed for the non-reachability maze experiment over 100 episodes.
@@
Experiment 2: Reachability (maze). As the second experiment, we consider the case where the primary objective is a reachability objective and the secondary objective is non-terminating, which was the setting that we found that TLQ fails to reach the goal state in Section 4.1. As the agent only cares about eventually reaching the goal, it can completely avoid going on a bad tile. We run Algorithm 3 for $N_{e}=4000$ episodes. Out of 10 seeds, 4 find policies that have 90% success over 100 episodes. Figure 5 shows how the satisfaction frequency for each objective changes throughout training.
Figure 5: Satisfaction rates for a single successful seed for the reachability maze experiment over 100 episodes.
@@
Experiment 3: Reachability (FTN). Finally, we present our results on Fruit Tree Navigation (FTN) [26]. This task requires the agent to explore a full binary tree of depth $d=5$ with fruits on the leaf nodes. Each fruit has a randomly assigned vectorial reward $\mathbf{r}\in\mathbb{R}^{6}$ which encodes the amount of different nutrition components of the fruit. The agent needs to find a path from the root to the fruit that fits to the user preferences by choosing between left and right subtrees at every non-leaf node. As it has non-zero rewards only at the leafs, this is a reachability objective.
This domain perfectly highlights the benefits of thresholded lexicographic user preference compared to linear scalarization. The user can have a certain threshold that needs to be reached for each nutrition component and an importance order between these components that should be followed if it is not possible to satisfy all of the thresholds. Using linear scalarization in this case requires knowing the reward values of all of the fruits beforehand, deciding which fruit would fit the user preference, and finding a weight vector $\omega$ for which the desired fruit is better than others. Figure 6 compares our algorithm with TLQ on this domain. It also shows the need for the $\Delta$ parameter and highlights that using just hyperplanes as done in [20] fails this task. Similarly, TLQ agent also fails to find the desired leaf.

### v2

@@ Abstract
Lexicographic multi-objective problems, which impose a lexicographic importance order over the objectives, arise in many real-life scenarios. Existing Reinforcement Learning work directly addressing lexicographic tasks has been scarce. The few proposed approaches were all noted to be heuristics without theoretical guarantees as the Bellman equation is not applicable to them. Additionally, the practical applicability of these prior approaches also suffers from various issues such as not being able to reach the goal state. While some of these issues have been known before, in this work we investigate further shortcomings and propose fixes for improving practical performance in many cases. We also present a policy optimization approach using our Lexicographic Projection Algorithm (LPA) that has the potential to address these theoretical and practical concerns. Finally, we demonstrate our proposed algorithms on benchmark problems.
@@ 1 Introduction
Our present work is in the a priori setting. The most common specification method in this setting is linear scalarization which which requires the designer to assign weights to the objectives and take a weighted sum, thus making solutions comparable [4]. The main benefit of this technique is that it allows the use of many standard off-the-shelf algorithms as it preserves the additivity of the reward functions. However, expressing user preference with this technique requires significant domain knowledge and preliminary work in most scenarios [14]. While it can be the preferred method when the objectives can be expressed in comparable quantities, e.g. when all objectives have a monetary value, this is not the case most of the time. Usually, the objectives are expressed in incomparable quantities like money, time, and carbon emissions. Additionally, a composite utility combining the various objectives, and an approximation of that with linear scalarization limits us to a subset of the Pareto optimal set.
To address these drawbacks of linear scalarization, several other approaches have been proposed and studied. Nonlinear scalarization methods like Chebyshev [20] are more expressive and can capture all of the solutions in the Pareto optimal set, however, they do not provide ease of writing. In this paper, we will focus on an alternative specification method that overcomes both limitations of linear scalarization, named Thresholded Lexicographic Ordering (TLO) [5] [14]. In lexicographic ordering, the user determines an importance order for the objectives. The more important objectives dominate the lower order objectives. Given two candidate solutions, a lower ranked objective is used to rank the two solutions only if all higher order objectives have the same values. The thresholding part of the technique allows a more generalized definition for being the same w.r.t. an objective. The user provides a threshold for each objective except the last, and the objective values are clipped at the corresponding thresholds. This allows the user to specify values beyond which they are indifferent to the optimization of an objective. There is no threshold for the last objective as it is considered an unconstrained open-ended objective.
@@ 2 Related Work
[5] was one of the first papers to investigate RL in multi-objective tasks with preference ordering. It introduces TLQ as an RL algorithm to solve such problems. Vamplew et al. [31] showed that TLQ significantly outperforms Linear Scalarization (LS) when the Pareto front is globally concave or when most of the solutions lie on the concave parts. However, LS performs better when the rewards are not restricted to terminal states, because TLQ cannot account for the already received rewards. Later, Roijers et al. [22] generalized this analysis by comparing more approaches using a unifying framework. To our knowledge, [31] is the only previous work that explicitly discussed shortcomings of TLQ. However, we found that TLQ has other significant issues that occur even outside of the problematic cases they analyze.
Wray et al. [35] introduced slack-based Lexicographic MDP (LMDP) and the Lexicographic Value Iteration (LVI) algorithm. These LMDPs specifies slacks from optimal policy instead of absolute thresholds. In other words, the user specifies how much worse than the optimal an objective can be instead of simply how good it should be. For example, in slack setting, one would say: “I need to get there at most 20% slower than the fastest I could”. In the thresholds setting, one would say: “I need to get there in 30 minutes”. These two settings require special methods. While Wray et al. [35] proved the convergence to desired policy if slacks are chosen appropriately, such slacks are generally too tight to allow user preferences. This is also observed in Pineda et al. [21] which claimed that while ignoring these slack bounds negates the theoretical guarantees, the resulting algorithm still can be useful in practice.
@@
However, we discovered that these works on using a Q-learning variant with thresholded ordering perform very poorly in most cases due to non-Markovianity of the value function they try to learn. It is possible to bypass this issue by using policy gradient approaches as they do not require learning an optimal value function. In order to handle conflicting gradients, some modifications to the gradient descent algorithm are needed. Recent work on modified gradient descent algorithms came mostly from Multi Task Learning literature, which could be considered a multiobjective optimization problem [3] [24] [16] [18] [17][10]. While these papers use similar ideas with our work, their setting is different than ours as they do not have any explicit importance order; hence, not applicable to our setting. Uchibe and Doya [30] have the most similar setting to ours in gradient-based algorithms. It considers a set of constraints with an unconstrained objective. Then, the gradient of the unconstrained objective is projected onto the positive half-space of the active (violated) constraints and adds a correction step to improve the active constraints. When no valid projection is found, the most violated constraints are ignored until a valid projection exists. This is one of the main differences with our setting: As we have an explicit importance-order of the objectives, it is not acceptable to ignore a constraint without considering the importance order. Also, we project the gradients onto hypercones instead of hyperplanes, which is a hypercone with $\pi/2$ vertex angle. Thus, our algorithm allows varying degrees of conservative projections to prevent a decline in the constraints.
While there are many other recent works on Constrained Markov Decision Process (CMDPs) like [33, 11], their approaches are not directly applicable as an importance order over constraints is not allowed. However, an LMDP problem can be reduced to a series of CMDP problems using the following observation. When all constraints can be satisfied simultaneously, the importance order between them becomes irrelevant for the optimal policy. Hence, the LMDP can be solved as a CMDP. Then, for a general LMDP problem where all constraints cannot be necessarily satisfied simultaneously, the main task is to find the maximum $k$ such that most important $k$ objectives can be satisfied together. The most straightforward way to achieve this would be starting by checking the feasibility of the most important constraint only. Then, we can keep adding the next constraint in the importance order until the problem becomes infeasible in order to find the maximum $k$. Note that this algorithm is similar to "linear search" and would require $O(K)$ CMDPs to be solved for an LMDP with $K$ objectives. One can implement different strategies that imitate other search algorithms like binary search.
@@
Skalse et al. [26] propose both value-based and policy-based approaches for LMDPs. Their value-based approach is based on slacks like [14] and they require using very small slack values. This protects their approach from the issues with having relaxations by limiting their setting to strict lexicographic order, i.e. not allowing any suboptimality for the more important objectives. For policy-based methods, they use Lagrangian relaxation and their setting is again a strict lexicographic ordering, i.e. it does not allow treating values above a threshold equal. As we are primarily interested in solving tasks with thresholded objectives, this paper does not address our question.
Finally, using RL with lexicographic ordering recently began to attract attention from other communities as well. For example, Hahn et al. [7] use formal methods to construct single objective MDPs when all of the objectives are $\omega$-regular.
@@ 4.1 Shortcomings of Prior TLQ Approaches
Endpoint Objective: An endpoint objective is represented by a non-zero reward in the terminal states of the MDP and zero rewards elsewhere. We will call an objective that has non-zero rewards in at least one non-terminal state a path objective.
@@
(I) TLQ successful case studies: Vamplew et al. [31] show an experiment where the constrained objective is endpoint, unconstrained objective is terminating, and TLQ works. Li and Czarnecki [14] show this for a case study where the constrained objective is a non-terminating endpoint constraint. Note that these are just demonstrations of empirical performance for case studies that fall into these categories and the papers do not make claims about general instances.
(II) TLQ does not work: Vamplew et al. [31] show that TLQ does not work when constrained objective is a path objective. To the best of our knowledge, this is the only failure case that has been shown previously in the literature.
@@
In addition to previously identified failure case in Vamplew et al. [31], TLQ does not work when the constrained objective is a terminating endpoint objective but the unconstrained one is non-terminating.
@@
Problems with Endpoint Constraint. A common scenario in control tasks is having a primary objective to reach a goal state and a secondary objective that evaluates the path taken to there. Formally, this is a scenario where the primary objective is a endpoint objective. However, TLQ either needs to ignore the secondary objective or fails to satisfy the primary objective in such cases. All of the thresholding methods above fail to guarantee to reach the goal in this setting when used threshold/slack values are uniform throughout state space. The maze in Figure 1 can be used to observe this phenomenon. Assume that our primary objective is to reach $G$ and we encode this with a reward function that is 00 everywhere but $G$ where it is $R$. And our secondary objective is to avoid the bad tiles. A Pareto optimal policy in this maze could be indicated by the state trajectory $(1,0)\to(2,0)\to(2,1)\to(2,2)\to(1,2)$. However, this is unattainable by TLQ.
Since the reward is given only in the goal state, $\hat{Q}^{\star}_{1}$ can be equal to $\tau_{1}$ only for the state-action pairs that lead to the goal state. All others will be discounted from these; hence, less than the threshold $\tau_{1}$. This means the agent will be focusing only on primary objective and ignore the secondary objective when using absolute thresholding of [5]. If no discounting was used, all actions would have the value $\tau_{1}$, hence the agent would not need to reach the goal state. We believe the reason why Vamplew et al. [31] have not observed this issue in their experiments with undiscounted ($\gamma=1$) Deep Sea Treasure (DST) is due to their objectives. The secondary objective of DST, minimizing the time steps before terminal state, is a terminating objective which pushes the agent to actually reach a goal state. Absolute Slacking also similarly leads to contradicting requirements when the trajectory described above is aimed but we leave this case to the reader due to space constraints.
@@ 5 Policy Gradient Approach for TLO
The proof and the exact lower bounds on $F_{j}(\boldsymbol{x}^{(l+1)})-F_{j}(\boldsymbol{x}^{(l)})$ in the proposition can be found in Appendix G.
@@
Experiment. As a benchmark for the Lexicographic Projection Algorithm, we used the objective functions $F_{1}(x,y)=-4x^{2}+-y^{2}+xy$ and $F_{2}(x,y)=-(x-1)^{2}-(y-0.5)^{2}$ which are taken from [37]. We modified $F_{2}$ slightly for better visualization and multiplied both functions with $-1$ to convert this to a maximization problem. We set the threshold for $F_{1}$ to $-0.5$. The behavior of our cone algorithm without using active constraints heuristic on this problem with $\tau=(-0.5)$ and $\Delta=\frac{\pi}{90}$ can be seen in Figure 2.
@@ 6 Experiments
We evaluate the performance of the Lexicographic REINFORCE algorithm on two Maze problems and a many-objective benchmark from the literature. In both experiments, we use a two layer neural network for policy function. See Appendix H.2.1 for further details on the policy function.
@@
Experiment 1: Path Objective. The primary objective is a path objective, i.e. it takes non-zero values in some non-terminal states. For this, we flip our objectives from the previous maze setting and define our primary objective as minimizing the cost incurred from the bad tiles. $HH$s give $-5$ reward and $hh$s give $-4$ reward. $+1$ reward is given in the terminal state to extend the primary objective to have rewards in both terminal and non-terminal states. The secondary objective is to minimize the time taken to the terminal state. We formalize this by defining our secondary reward function as 00 in terminal state and $-1$ everywhere else.We use the Maze in Figure 3 for this experiment.
Figure 4: Satisfaction rates for a single successful seed for the path objective maze experiment over 100 episodes.
@@
Experiment 2: Endpoint objective (maze). As the second experiment, we consider the case where the primary objective is an endpoint objective and the secondary objective is non-terminating, which was the setting that we found that TLQ fails to reach the goal state in Section 4.1. As the agent only cares about eventually reaching the goal, it can completely avoid going on a bad tile. We run Algorithm 3 for $N_{e}=4000$ episodes. Out of 10 seeds, 4 find policies that have 90% success over 100 episodes. Figure 5 shows how the satisfaction frequency for each objective changes throughout training. Further experimental details and results can be found in Appendix H.2.2.
Figure 5: Satisfaction rates for a single successful seed for the endpoint maze experiment over 100 episodes.
@@
Experiment 3: Endpoint Objective (FTN). Finally, we present our results on Fruit Tree Navigation (FTN) [36]. This task requires the agent to explore a full binary tree of depth $d=5$ with fruits on the leaf nodes. Each fruit has a randomly assigned vectorial reward $\mathbf{r}\in\mathbb{R}^{6}$ which encodes the amount of different nutrition components of the fruit. The agent needs to find a path from the root to the fruit that fits to the user preferences by choosing between left and right subtrees at every non-leaf node. As it has non-zero rewards only at the leafs, this is an endpoint objective.
Figure 6 compares our algorithm with TLQ on this domain. It also shows the need for the $\Delta$ parameter and highlights that using just hyperplanes as done in [30] fails this task. Similarly, TLQ agent also fails to find the desired leaf.
@@
6.1 Discussions and Comparison with Other Methods
In this section, we will try to summarize the significance of our results w.r.t. other methods from the literature. The first two experiments showcase our method’s effectiveness in settings where TLQ is shown to fail. Since TLQ completely fails in these settings, it is omitted from the plots. In Experiment 1, the constrained objective is a path objective which is the setting that Vamplew et al. [31] has found that TLQ fails. However, our experiments show that Lexicographic REINFORCE can actually solve this setting. Similarly, Experiment 2 deals with the setting where the constrained objective is a terminating endpoint objective and the unconstrained objective is non-terminating, a setting we show to be unsolvable by TLQ in Section 4.1. Our results show that Lexicographic REINFORCE can solve this problem too.
While Experiment 1 and 2 show that our method does not suffer from the inherent limitations observed in TLQ, they do not show how our method compares to TLQ in a setting that is not picked to highlight the shortcomings of TLQ. For this reason, we test our method on FTN benchmark from the literature. This benchmark has endpoint constraints and the unconstrained objective is terminating, a setting which was shown to be solvable by TLQ ([31]). However, our results in Figure 6 suggest that as the number of objectives increases, TLQ may fail to find an optimal policy. In contrast, Lexicographic REINFORCE can solve this task with the right $\Delta$ parameter. The results also suggest that the hyperplane projection method of [30] is not sufficient and the conservativeness of hypercone projection may be needed to successfully converge to the optimal policy in a challenging task. Note that while the performance of Lexicographic REINFORCE is sensitive to $\Delta$ parameter, we believe that combining our method with a more stable policy-gradient algorithm could reduce this sensitivity significantly.
Note that we believe the baselines used in the FTN experiment are the only methods that are directly comparable with our algorithm. In particular, we believe linear scalarization does not offer a fair comparison with our method as it puts a great burden on the user to specify their preferences. Using linear scalarization in this case requires knowing the reward values of all of the fruits beforehand, deciding which fruit would fit the user preference, and finding a weight vector $\omega$ for which the desired fruit is better than others. (Note that such an $\omega$ may not even exist if the desired fruit lies in the concave part of the Pareto front.) In contrast, our method requires only the importance order and the threshold values. However, we also note that if finding such an $\omega$ is not an issue, linear scalarization allows using any method from the vast single-objective RL literature and can yield a great learning performance.